# latexText page 4 contains only floats

### v2

@@ Abstract
Timely and accurate assessments of building damage are crucial for effective response and recovery in the aftermath of earthquakes. Conventional preliminary damage assessments (PDA) often rely on manual door-to-door inspections, which are not only time-consuming but also pose significant safety risks. To safely expedite the PDA process, researchers have studied the applicability of satellite imagery processed with heuristic and machine learning approaches. These approaches output binary or, more recently, multiclass damage states at the scale of a block or a single building. However, the current performance of such approaches limits practical applicability. To address this limitation, we introduce a metadata-enriched, transformer-based framework that combines high-resolution post-earthquake satellite imagery with building-specific metadata relevant to the seismic performance of the structure. Our model achieves state-of-the-art performance in multiclass post-earthquake damage identification for buildings from the Mw 7.8 Turkey-Syria earthquake on February 6, 2023. Specifically, we demonstrate that incorporating metadata, such as seismic intensity indicators, soil properties, and synthetic aperture radar (SAR) damage proxy maps not only enhances the model’s accuracy and ability to distinguish between damage classes, but also improves its generalizability across various regions. Furthermore, we conducted a detailed, class-wise analysis of feature importance to understand the model’s decision-making across different levels of building damage. This analysis reveals how individual metadata features uniquely contribute to predictions for each damage class. By leveraging both satellite imagery and metadata, our proposed framework enables faster and more accurate damage assessments for precise, multiclass, building-level evaluations that can improve disaster response and accelerate recovery efforts for affected communities.
@@ 1 Introduction
In the aftermath of an earthquake, conducting a Preliminary Damage Assessment (PDA) is typically the first step in evaluating the extent of damage and determining the need for governmental assistance FEMA (Accessed: 2023-06-05). The PDA process typically begins at the local level, where initial damage data is collected by local authorities through door-to-door inspections Reliefweb (2023) to determine the magnitude of damage, the impact of the disaster, and identify the worst-hit regions. These inspections can be unsafe due to falling hazards and other safety concerns due to loss of structural integrity or the risk of aftershocks. Additionally, delays from the large number of structures to be assessed can also significantly hinder the speed of governmental aid deployment, exacerbating the recovery period for affected communities. In 2017, following the Mexico City earthquake, the Civil Engineering Association of Mexico took approximately three weeks to complete evaluations across the entire city, which delayed recovery efforts and left thousands of people without access to their homes Wang et al. (2022a); usnews (Accessed: 2023-06-05). Similarly, in 2023, the Turkey–Syria earthquakes affected 15 million people, with the preliminary damage assessments taking more than one month to complete. Erdemir et al. (2023); Koca and Topak (2024); Okuyama et al. (2024); Tena-Colunga et al. (2021); preventionweb (Accessed: 2023-06-05). These cases underscore a critical need for alternative methods that can accelerate the PDA process for quicker and more effective response effort.
@@
The time-consuming part of the PDA process is the door-to-door building damage assessment. To help speed up access to information for decision making, researchers have proposed frameworks to automate the post-earthquake damage assessment process for the identification of the damage state at the individual building level Foroughnia et al. (2024); Whitworth et al. (2022); Voelker et al. (2024). We discuss these frameworks by adopting a categorization based on two criteria: (i) data type, including images from unmanned aerial vehicles (UAVs) and satellites and (ii) data processing methods, either heuristic-based approaches relying on predefined rules or data-driven approaches, such as machine learning (ML) and deep learning (DL) techniques, which learn patterns and make predictions based on past data.
@@
Researchers have employed various data sources for post-earthquake damage assessment, including aerial imagery obtained from UAVs Li et al. (2022); Liu et al. (2024); Zhang et al. (2023), synthetic aperture radar (SAR) satellite data Ge et al. (2020); Macchiarulo et al. (2023); Bai et al. (2017); Rao et al. (2023); Brunner et al. (2010); Macchiarulo et al. (2024); Yang et al. (2024), and optical satellite imagery Ahmadi et al. (2023); Xia et al. (2023a); Wang et al. (2022b); Joshi et al. (2021); Wang et al. (2023); Ilmak et al. (2024). Data collection with UAVs can result in high-resolution and actionable imagery as demonstrated in Liu et al. (2024); Zhang et al. (2023). However, scaling UAV data collection to large areas can be time consuming and challenging due to restrictions on flying beyond visual line of sight (BVLOS) and limited battery life. For earthquakes that affect large areas, satellite imagery can be very valuable due to its wide coverage, time efficiency, and on-demand availability. In a paper by Rao et al. (2023) the authors attempted multiclass and binary damage classification for four recent earthquakes using SAR imagery. For three out of the four earthquakes studied, the trained model achieved close to 50% accuracy in detecting damaged buildings using binary classification . Similarly, Ahmadi et al. (2023) proposed a two-stage, dual-branch UNet architecture with shared weights between the two branches using optical data . However, the paper highlights the problem of low accuracy for intermediate classes in multiclass classification. Macchiarulo et al. utilized post-earthquake SAR imagery with a method that involved combining textural features from the imagery. The output was a map that classifies damage into five damage states at the level of city blocks, each consisting of an average of 48 buildings. However, this approach is not suitable for building-level damage assessment Macchiarulo et al. (2023); Giardina et al. (2024). A common limitation in these papers is the challenge of multiclass building-level damage assessment, primarily because nadir images make it difficult to differentiate between intermediate damage classes in earthquake-affected buildings. Additionally, the resolution of SAR imagery is often insufficient for precise building-level predictions.
@@
On the basis of data processing methods, frameworks can be categorized into heuristic-based Arciniegas et al. (2007); Yun et al. (2015); Fielding and Jung (2024); Loos et al. (2020) and data-driven approaches Ahmadi et al. (2023); Bai et al. (2017); Joshi et al. (2021); Zhang et al. (2023); Tamkuan and Nagai (2017); Kalantar et al. (2020); Xia et al. (2023a); Varghese and Hoskere (2023); Genova et al. (2024); Ghimire et al. (2022). Heuristic-based methods, such as the normalized difference built-up index (NDBI) derived from optical satellite imagery, utilize electromagnetic radiation to detect surface changes in objects like vegetation and building Yu et al. (2024). While NDBI effectively highlights impervious surfaces such as buildings, it is susceptible to cloud occlusion. In contrast, SAR-derived indices like the amplitude dispersion index (ADI) Esmaeili and Motagh (2016) and damage proxy maps (DPMs) Yun et al. (2015) are not affected by environmental factors ESA (Accessed: 2023-06-05). DPMs derived from bi-temporal SAR imagery, are frequently used for rapid post-earthquake damage assessment. However, due to low resolution, DPMs are less effective at the building level, as seen in the weak correlation with building damage during the 2015 Nepal earthquake Loos et al. (2020); Yun et al. (2015). Furthermore, damage maps produced using different satellite-based remote sensing techniques often show significant discrepancies, and extensive validation data are still required to accurately characterize the performance of these methods at both high and medium resolutions Voelker et al. (2024).
@@
Data-driven methods Ahmadi et al. (2023); Bai et al. (2017); Joshi et al. (2021); Zhang et al. (2023), which include machine learning (e.g., support vector machines (SVM), Random Forest (RF), etc.) and deep learning approaches (e.g., U-net, ResNet, etc.), are increasingly being applied for post-earthquake assessments. ML models are generally trained on tabular data. Rao et al. (2023) proposed a ML framework that utilizes earthquake-related indices, such as peak ground acceleration (PGA), and DPMs for building-level damage assessment. Similarly, Yu et al. further combined NDBI and ADI with DPMs and PGA to train a random forest model for PDA Yu et al. (2024). The result shows ML methods outperform the traditional practice of relying solely on DPMs. However, the results are still limited to binary classification and do not perform well in multiclass classification scenarios. On the other hand, the training of deep learning approaches often require large amounts of high-resolution optical imagery. Recent studies Ahmadi et al. (2023); Xia et al. (2023b) have proposed convolutional neural network-based two-staged models for localization and classification. The results clearly demonstrate that deep neural network architectures outperform SVM and Random Forest in binary building damage classification. However, low accuracy in multiclass classification remains a challenge. One major reason for the low accuracy in multiclass classification is the imbalanced dataset. For instance, in the largest disaster dataset, xBD Gupta et al. (2019), which categorizes building damage into four levels, only 3 out of 51,500 buildings are classified as completely destroyed, highlighting the severe class imbalance.
These studies highlight that each data type and methodology offers unique insights, along with certain limitations. For instance, earthquake-related indices derived from SAR provide valuable insights into structural changes, PGA offers local estimates of seismic intensity of the event, and optical imagery offers detailed visual information. However, the current state-of-the-art methods don’t fully utilize all of these sources and primarily rely on satellite imagery. As a result, they fall short of achieving the required accuracy for practical use, particularly in multiclass classification scenarios. Additionally, Singh et al. demonstrated a significant improvement in post-hurricane building assessments using transformer vs CNN models, motivating the study of these methods for earthquakes Singh and Hoskere (2023a, b). This underscores the importance of carefully integrating these different approaches to develop a more effective post-disaster assessment (PDA) framework.
@@
This study proposes a novel PDA framework for earthquakes (Figure 1) integrating high-resolution post-earthquake satellite imagery with publicly available metadata with a novel transformer-based network. Metadata includes geographical attributes, disaster intensity variables, and soil properties. The novel contributions of our research include (i) the development of an earthquake-specific multiclass dataset at the building level with corresponding metadata, (ii) a new deep learning approach termed QuakeMetaFormer, a transformer-based model that integrates metadata with optical imagery for post-earthquake PDA, and (iii) a class-wise comparative analysis of the impact of different metadata on damage assessment. This paper is organized as follows. Section 2 provides a detailed discussion of the proposed framework, including the processes of data collection and processing, and the architecture of QuakeMetaFormer. Section 3 describes four experiments conducted to test our hypotheses: (i) the impact of adding metadata on model accuracy, (ii) comparison of QuakeMetaFormer and previously published machine learning frameworks, (iii) feature importance across various metadata and (iv) testing generalization capability of QuakeMetaFormer across regions. Section 4 presents and analyzes the results from these experiments. Finally, Section 5 concludes the paper by summarizing the key findings and their implications.
@@ 2.1 Dataset Collection
To evaluate our proposed framework, we curate a novel dataset from buildings affected by the Turkey-Syria earthquake of February 6, 2023, where a magnitude 7.8 quake was centered near Kahramanmaraş in southern Turkey, as shown in Figure 2(a). Strong aftershocks, including those of magnitudes 7.5 and 7.6 ReliefWeb (Accessed: 2023-03-05), further contributed to the widespread destruction, causing additional damage. The earthquake damaged more than 200,000 buildings and led to the collapse of approximately 20,000 structures. Erdemir et al. (2023); Okuyama et al. (2024). We now discuss details involved in the extraction of satellite imagery and corresponding metadata.
Figure 2: (a) Map showing areas of the study affected by the Turkey-Syria earthquake, and (b) post-earthquake satellite imagery provided by MAXAR Technologies, with corresponding building footprints obtained from OpenStreetMap.
@@ 2.1.1 Satellite Imagery
The satellite imagery used in this study were obtained through MAXAR Technologies’ open data program Maxar (Accessed: 2023-06-05b), which provides before and after satellite images during disasters. These images are geotagged and have a resolution of 30 cm ground sample distance (GSD) Maxar (Accessed: 2023-06-05a). In this study, we exclusively used post-earthquake imagery and filtered the images based on the areas affected by the earthquake. To extract building-level images, corresponding building footprints from OpenStreetMap (OSM) were utilized OSM (Accessed: 2023-06-05). OSM is a free and open-source geographic database, that is continuously updated and maintained by a global community. Figure 2 illustrates a sample from the dataset for the city Kahramanmaraş, highlighting the satellite imagery and corresponding building footprints. The final output for each building is a cropped section of the satellite imagery, containing three-channel (RGB) pixel-level data within the boundaries of the building footprint.

### v3

@@ Abstract
Timely and accurate assessments of building damage are crucial for effective response and recovery in the aftermath of earthquakes. Conventional preliminary damage assessments (PDA) often rely on manual door-to-door inspections, which are not only time-consuming but also pose significant safety risks. To safely expedite the PDA process, researchers have studied the applicability of satellite imagery processed with heuristic and machine learning approaches. These approaches output binary or, more recently, multiclass damage states at the scale of a block or a single building. However, the current performance of such approaches limits practical applicability. To address this limitation, we introduce a metadata-enriched, transformer-based framework that combines high-resolution post-earthquake satellite imagery with building-specific metadata relevant to the seismic performance of the structure. Our model achieves state-of-the-art performance in multiclass post-earthquake damage identification for buildings from the Mw 7.8 Türkiye-Syria earthquake on February 6, 2023. Specifically, we demonstrate that incorporating metadata, such as seismic intensity indicators, soil properties, and synthetic aperture radar (SAR) damage proxy maps not only enhances the model’s accuracy and ability to distinguish between damage classes, but also improves its generalizability across different regions affected by an earthquake event. Furthermore, we conducted a detailed, class-wise analysis of feature importance to understand the model’s decision-making across different levels of building damage. This analysis reveals how individual metadata features uniquely contribute to predictions for each damage class. By leveraging both satellite imagery and metadata, our proposed framework enables faster and more accurate damage assessments for precise, multiclass, building-level evaluations that can improve disaster response and accelerate recovery efforts for affected communities.
@@ 1 Introduction
In the aftermath of an earthquake, conducting a Preliminary Damage Assessment (PDA) is typically the first step in evaluating the extent of damage and determining the need for governmental assistance FEMA (Accessed: 2023-06-05). The PDA process typically begins at the local level, where initial damage data is collected by local authorities through door-to-door inspections Reliefweb (2023) to determine the magnitude of damage, the impact of the disaster, and identify the worst-hit regions. These inspections can be unsafe due to falling hazards and other safety concerns due to loss of structural integrity or the risk of aftershocks. Additionally, delays from the large number of structures to be assessed can also significantly hinder the speed of governmental aid deployment, exacerbating the recovery period for affected communities. In 2017, following the Mexico City earthquake, the Civil Engineering Association of Mexico took approximately three weeks to complete evaluations across the entire city, which delayed recovery efforts and left thousands of people without access to their homes Wang et al. (2022a); usnews (Accessed: 2023-06-05). Similarly, in 2023, the Türkiye–Syria earthquakes affected 15 million people, with the preliminary damage assessments taking more than one month to complete. Erdemir et al. (2023); Koca and Topak (2024); Okuyama et al. (2024); Tena-Colunga et al. (2021); preventionweb (Accessed: 2023-06-05). These cases underscore a critical need for alternative methods that can accelerate the PDA process for quicker and more effective response efforts.
@@
The time-consuming part of the PDA process is the door-to-door building damage assessment. To help speed up access to information for decision making, researchers have proposed frameworks to automate the post-earthquake damage assessment process for the identification of the damage state at the individual building level Foroughnia et al. (2024); Whitworth et al. (2022); Voelker et al. (2024). We discuss these frameworks by adopting a categorization based on two criteria: (i) data type, including images from unmanned aerial vehicles (UAVs) and satellites and (ii) data processing methods, either heuristic-based approaches relying on predefined rules (e.g., Arciniegas et al. (2007); Yun et al. (2015); Fielding and Jung (2024); Loos et al. (2020)) or data-driven approaches, such as machine learning (ML) and deep learning (DL) techniques, which learn patterns and make predictions based on past data and data-driven approaches (e.g., Ahmadi et al. (2023); Bai et al. (2017); Joshi et al. (2021); Zhang et al. (2023); Tamkuan and Nagai (2017); Kalantar et al. (2020); Xia et al. (2023a); Varghese and Hoskere (2023); Genova et al. (2024); Ghimire et al. (2022)).
@@
Researchers have employed various data sources for post-earthquake damage assessment, including aerial imagery obtained from UAVs Li et al. (2022); Liu et al. (2024); Zhang et al. (2023), synthetic aperture radar (SAR) satellite data Ge et al. (2020); Macchiarulo et al. (2023); Bai et al. (2017); Rao et al. (2023); Brunner et al. (2010); Macchiarulo et al. (2024); Yang et al. (2024), and optical satellite imagery Ahmadi et al. (2023); Xia et al. (2023a); Wang et al. (2022b); Joshi et al. (2021); Wang et al. (2023); Ilmak et al. (2024). Data collection with UAVs can result in high-resolution and actionable imagery as demonstrated in Liu et al. (2024); Zhang et al. (2023). However, scaling UAV data collection to large areas can be time consuming and challenging due to restrictions on flying beyond visual line of sight (BVLOS) and limited battery life. For earthquakes that affect large areas, satellite imagery can be very valuable due to its wide coverage, time efficiency, and on-demand availability. In a paper by Rao et al. (2023) the authors attempted multiclass and binary damage classification for four recent earthquakes using SAR imagery. For three out of the four earthquakes studied, the trained model achieved close to 50% accuracy in detecting damaged buildings using binary classification. Similarly, Ahmadi et al. (2023) proposed a two-stage, dual-branch UNet architecture with shared weights between the two branches using bi-temporal optical data. However, the paper highlights the problem of low accuracy for intermediate classes in multiclass classification. Additionally, a common limitation of change detection algorithms for PDA is that they rely on bi-temporal data. However, in many cases, pre-disaster images may not be readily available near the disaster date, or the built environment may have undergone significant changes since the last image was taken Vitale and Milillo (2024). Macchiarulo et al. (2023) and Giardina et al. (2024) utilized post-earthquake SAR imagery with a method that involved combining textural features from the imagery. The output was a map that classifies damage into five damage states at the level of city blocks, each consisting of an average of 48 buildings. However, this approach is not suitable for building-level damage assessment. A common limitation in these papers is the challenge of multiclass building-level damage assessment, primarily because nadir images make it difficult to differentiate between intermediate damage classes in earthquake-affected buildings. Additionally, the resolution of SAR imagery is often insufficient for precise building-level predictions.
@@
Heuristic-based methods, such as the normalized difference built-up index (NDBI) derived from optical satellite imagery, utilize electromagnetic radiation to detect surface changes in objects like vegetation and buildings Yu et al. (2024). While NDBI effectively highlights impervious surfaces such as buildings, it is susceptible to cloud occlusion. In contrast, SAR-derived indices like the amplitude dispersion index (ADI) Esmaeili and Motagh (2016) and damage proxy maps (DPMs) Yun et al. (2015) are not affected by environmental factors ESA (Accessed: 2023-06-05). DPMs derived from bi-temporal SAR imagery are frequently used for rapid post-earthquake damage assessment. However, due to low resolution, DPMs are less effective at the building level, as seen in the weak correlation with building damage during the 2015 Nepal earthquake Loos et al. (2020); Yun et al. (2015). Furthermore, damage maps produced using different satellite-based remote sensing techniques often show significant discrepancies, and extensive validation data are still required to accurately characterize the performance of these methods at both high and medium resolutions Voelker et al. (2024).
@@
Data-driven approaches including machine learning (e.g., support vector machines (SVM), Random Forest (RF), etc.) and deep learning (e.g., U-net, ResNet, etc.), have widely been adopted for modern post-earthquake assessments. ML models are generally trained on tabular data. Rao et al. (2023) proposed a ML framework that utilizes earthquake-related indices, such as peak ground acceleration (PGA), and DPMs for building-level damage assessment. Yu et al. (2024) further combined NDBI and ADI with DPMs and PGA to train a random forest model for PDA. The result shows ML methods outperform the traditional practice of relying solely on DPMs. However, the results are still limited to binary classification and do not perform well in multiclass classification scenarios. Li et al. (2025); Xu et al. (2022) introduced a variational causal Bayesian network to assess earthquake-induced building damage, landslides, and liquefaction using 30m bi-temporal SAR imagery, USGS shakemaps, and building footprints, but without optical data. The model outputs a 30m-resolution raster with probability estimates (ranging from 0 to 1) for each pixel. The result shows improved class discrimination compared to using only DPMs. However, the authors state that these probabilities are event-dependent e.g., both low- and high-magnitude events yield values between 0 and 1, making it difficult to compare damage severity across events and locations. The generalization capability of the approach for classifying building damage to new events is thus limited by the need for multiple thresholds to be tuned based on the ground truth acquired by other means for the event. Additionally, when multiple buildings fall within a single pixel (likely to occur in 30m resolution imagery), they all receive the same probability, making it difficult to disambiguate damage classes for a single building. Given that modern electro-optical satellite imagery is available at a 100x better resolution of 0.3m, a significant research gap in improving the resolution of assessments by an order of magnitude would be analyzing the utility of optical imagery in providing a generalizable multiclass building-level assessment without the need for manual threshold tuning. On the other hand, the training of deep learning approaches often require large amounts of high-resolution optical imagery. Recent studies Ahmadi et al. (2023); Xia et al. (2023b) have proposed convolutional neural network-based two-staged models for localization and classification. The results clearly demonstrate that deep neural network architectures outperform SVM and Random Forest in binary building damage classification. However, low accuracy in multiclass classification remains a challenge. For multiclass classification, Singh et al. demonstrated a significant improvement in post-hurricane building assessments using transformer vs CNN models, motivating the study of these methods for earthquakes Singh and Hoskere (2023a, b).
Despite the advances described, practical implementations of post-disaster multiclass building PDA for earthquakes are challenging due to three main reasons. First, severe class imbalance in datasets remains a critical obstacle to achieving high accuracy in multiclass classification. For example, in the widely used large-scale bi-temporal satellite imagery dataset, xBD Gupta et al. (2019), which categorizes building damage into four severity levels, only 3 out of 51,500 buildings are labeled as completely destroyed. This extreme imbalance makes the training challenging and often leads to poor performance on minority classes. Second, generalization of damage assessment models across different regions and scenarios is limited as most statistical learning algorithms assume that data are independent and identically distributed (i.i.d.), meaning the training and test data come from the same underlying distribution. However, this assumption often fails in post-disaster scenarios due to regional variability, structural types, damage patterns, and limited data. Consequently, ensuring reliable model performance across diverse settings becomes challenging Quiñonero-Candela et al. (2022). Third, current approaches fail to integrate diverse data sources that offer complementary information for PDA. For instance, earthquake-related indices derived from SAR provide valuable insights about variations observed on the ground surface, PGA offers local estimates of seismic intensity of the event, and optical imagery offers detailed visual information. Yet, state-of-the-art methods predominantly focus either on satellite imagery, or the metadata like PGA, DPM, etc.. This limited data fusion restricts overall classification accuracy, particularly for multiclass damage assessments. Together, these challenges highlight the need for a carefully designed PDA framework that addresses data imbalance, improves generalizability, and effectively integrates various data sources aiming to enhance multiclass PDA at building level.
@@
This study proposes a novel multiclass building-level PDA framework for earthquakes (Figure 1) integrating high-resolution post-event satellite imagery with publicly available metadata with a novel transformer-based network. Metadata includes geographical attributes, disaster intensity variables, and soil properties. The novel contributions of our research include (i) the development of an earthquake-specific multiclass dataset at the building level with corresponding metadata, (ii) a transformer-based model termed QuakeMetaFormer that integrates metadata with high-resolution optical imagery for post-earthquake PDA while addressing severe class imbalance, (iii) a class-wise comparative analysis of the impact of different metadata on damage assessment, and (iv) an evaluation demonstrating the effectiveness of metadata in improving generalization across different regions affected by an earthquake. This paper is organized as follows. Section 2 provides a detailed discussion of the proposed framework, including the processes of data collection and processing, and the architecture of QuakeMetaFormer. Section 3 describes four experiments conducted to test our hypotheses: (i) the impact of adding metadata on model accuracy, (ii) comparison of QuakeMetaFormer and previously published machine learning frameworks, (iii) feature importance across various metadata and (iv) testing generalization capability of QuakeMetaFormer by integrating metadata. Section 4 presents and analyzes the results from these experiments. Finally, Section 5 concludes the paper by summarizing the key findings and their implications.
@@ 2.1 Dataset Collection
To evaluate our proposed framework, we curate a novel dataset from buildings affected by the Türkiye-Syria earthquake of February 6, 2023, where a magnitude 7.8 quake was centered near Kahramanmaraş in southern Türkiye, as shown in Figure 2(a). Strong aftershocks, including those of magnitudes 7.5 and 7.6 ReliefWeb (Accessed: 2023-03-05), further contributed to the widespread destruction, causing additional damage. The earthquake damaged more than 200,000 buildings and led to the collapse of approximately 20,000 structures. Erdemir et al. (2023); Okuyama et al. (2024). We now discuss details involved in the extraction of satellite imagery and corresponding metadata.
Figure 2: (a) Map showing areas of the study affected by the Türkiye-Syria earthquake, and (b) post-earthquake satellite imagery provided by MAXAR Technologies, with corresponding building footprints obtained from OpenStreetMap.
@@ 2.1.1 Satellite Imagery
The optical satellite imagery used in this study were obtained through MAXAR Technologies’ open data program Maxar (Accessed: 2023-06-05b), which provides before and after satellite images during disasters. These images are geotagged and have a resolution of 30 cm ground sample distance (GSD) Maxar (Accessed: 2023-06-05a). In this study, we exclusively used post-earthquake imagery and filtered the images based on the areas affected by the earthquake. To extract building-level images, corresponding building footprints from OpenStreetMap (OSM) were utilized OSM (Accessed: 2023-06-05). OSM is a free and open-source geographic database, that is continuously updated and maintained by a global community. Figure 2 illustrates a sample from the dataset for the city Kahramanmaraş, highlighting the satellite imagery and corresponding building footprints. The final output for each building is a cropped section of the satellite imagery, containing three-channel (RGB) pixel-level data within the boundaries of the building footprint.